OpenStudy (anonymous):

Find the matrix equation that represents the system of equations: 3x=5y+2 and y=15-6x

OpenStudy (mathteacher1729):

Start by putting each equation into the form: \[ax + by = c\]

OpenStudy (anonymous):

So: 3x-5y=2 -6x+y=15

OpenStudy (mathteacher1729):

Correct. Now the matrix form is: \[\begin{bmatrix}3 & 5 & 2 \\ -6 & 1 & 15\end{bmatrix}\]

OpenStudy (anonymous):

But how do I make the matrix equation out of that?

OpenStudy (mathteacher1729):

Oh, they want you to write: \[\begin{bmatrix}3&5\\ -6&1 \end{bmatrix}\begin{bmatrix}x\\y\end{bmatrix}=\begin{bmatrix}2\\15 \end{bmatrix}\]

OpenStudy (anonymous):

make a sl ut out of that..!

OpenStudy (mathteacher1729):

Do you mean row-reduce?

OpenStudy (anonymous):

Nope I got it now thank you so much!